Circular No. 6978

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998df IN ANONYMOUS GALAXY J. Maza, Department of Astronomy, University of Chile, reports the

discovery of an apparent supernova (B ∼ 18.5) by R. Antezana on a T-Max 400 film taken by L. E. González using the Maksutov telescope at Cerro El Roble on July 18.058 UT. SN 1998df is located at α = 21^{h}32^{m}30^{s}.57, δ = −65^{o}03′32′′.8 (equinox 2000.0), which is 1′′.7 west and 6′′.8 south of the nucleus of the host spiral galaxy. SN 1998df is also present on films obtained on July 1.175 (B ∼ 18.0) and June 1.350 (B ∼ 17.5). Several films obtained during 1997 show no object at the position of SN 1998df (B > 20).

SUPERNOVA 1998dg IN MCG –3-34-008 Maza also reports the discovery of an apparent supernova (B ∼ 17.5)

by Marina Wischnjewsky on a film taken (as above) on July 19.063 UT. SN 1998dg is located at α = 13^{h}10^{m}23^{s}.85, δ = −16^{o}56′08′′.1 (equinox 2000.0), which is 17′′.5 west and 17′′.1 south of the nucleus of MCG –3-34-008. The new star was confirmed on direct CCD images by E. Costa (University of Chile) on July 24.9 using the 0.91-m telescope at Cerro Tololo; a preliminary estimate of the magnitude is B ∼ 16. The object is not present on a film obtained on June 18.042 (B > 20).

SUPERNOVA 1998dh IN NGC 7541 W. D. Li, M. Modjaz, E. Halderson, T. Shefler, J. Y. King, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998dh was found and confirmed in unfiltered CCD images obtained on July 20.5 (mag ∼ 16.8) and 24.5 UT (mag ∼ 15.3). The new object is located at α = 23^{h}14^{m}40^{s}.31, δ = +4^{o}32′14′′.1 (equinox 2000.0), which is ∼ 53′′.5 west and 10′′.4 north of the nucleus of NGC 7541. A KAIT image obtained on July 16.5 did not show a star at the position of SN 1998dh (limiting mag ∼ 19.0).

SUPERNOVA 1998aq IN NGC 3982 Additional CCD V magnitude estimates by D. Hanžl, Brno, Czech

Republic (cf. IAUC 6937): June 17.955 UT, 14.23; 19.966, 14.4; 20.906, 14.4; 21.948, 14.3; 25.971, 14.4; 29.967, 14.45; July 20.977, 14.86.

1998 July 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green